Sam is 24 years younger than his mother. The total sum of their ages ( adding them together)is 72. How old is Sam's mother ?
Answer by haileytucki(390)

x+y=72_x=72-24
Subtract 24 from 72 to get 48.
x+y=72_x=48
Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is 48.
(48)+y=72_x=48
Remove the parentheses around the expression 48.
48+y=72_x=48
Since 48 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 48 from both sides.
y=-48+72_x=48
Add 72 to -48 to get 24.
y=24_x=48
Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is 24.
y=24_x=48